OpenStudy (anonymous):

PLease help. Divide 3x3 + 3x2 + 2x – 2 by x + 3 using long division.

OpenStudy (australopithecus):

watch this video https://www.youtube.com/watch?v=l6_ghhd7kwQ

OpenStudy (anonymous):

im still confused.. /;

OpenStudy (anonymous):

Do you still need help??

OpenStudy (anonymous):

yess

OpenStudy (anonymous):



OpenStudy (anonymous):

you are amazing, thank you so much! you made it clear too!

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

can you help me with one more?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

What are the real or imaginary solutions of the polynomial equation? 27x3 + 125 = 0

OpenStudy (anonymous):

27x^3+125=0 is the equation?

OpenStudy (anonymous):

yupp

OpenStudy (anonymous):

Do you also want the work?

OpenStudy (anonymous):

yes please, that would actually help! i understand it better

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

27x^3+125=0.... so factor the left side (3x+5)(9x^2-15x+25)=0 Now split each side (3x+5)=0 (9x^2-15x+25)=0 And solve each one separably. For the first one: (3x+5)=0 ... subtract 5 and then divide by 3 You get x=-5/3 Now solve the other part using the quadratic equation.

OpenStudy (anonymous):

Do you know how to use the quadratic formula? and do you follow so far?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

so i did the quadratoc formula and my answers i got

OpenStudy (anonymous):

and the other answer the same thing but its a minus sign between them am i right?

OpenStudy (anonymous):

I'll just show you what I got.

OpenStudy (anonymous):